# The Power of Linear Programming in Sponsored Listings Ranking: Evidence from a Large-Scale Field Experiment

### v3

@@ 1 Introduction
A widely adopted approach for the optimization stage in practice is the score-based ranking algorithm, which assigns each item a composite score based on predefined criteria and ranks items according to these scores. These algorithms are favored for their computational efficiency, typically producing rankings within 0.1 seconds to satisfy real-time latency requirements. The scoring criteria generally integrate multiple objectives, such as predicted consumer action rate, item relevance, and potential revenue generation. In most implementations, platforms employ a linear combination of these objectives to construct the score function that best aligns with their business priorities (Liu et al.,, 2015; Cheng et al.,, 2016; Covington et al.,, 2016; Okura et al.,, 2017; Zhou et al.,, 2018).
@@ 1.1 Literature reivew
Our work provides an approach for the ranking step in the recommendation system. In a recommendation system, the first step is to estimate the conversion rate of displaying the impression by machine learning algorithms, which has been quite mature and provide reasonable estimations in the online advertising industry (Liu et al.,, 2015; Cheng et al.,, 2016; Covington et al.,, 2016; Okura et al.,, 2017; Zhou et al.,, 2018). After estimating the conversion rates, online platforms are tasked with developing a ranking model that displays candidate items to consumers in a designated order, which is the focus of this paper. A majority of previous studies focus on the score-based ranking model (Zehlike et al.,, 2022; Renda and Straccia,, 2003; Montague and Aslam,, 2001; Manmatha et al.,, 2001), that is, calculating a score for each item through a customized scoring function, and then selecting the displayed items by ranking the scores. While the ranking algorithms are similar among various studies, the scoring function, however, might vary, depending on the objectives of each online platform. For instance, Covington et al., (2016) deployed a deep neural network model and used logistic regressions to generate scores for each video content on YouTube. Google Play (Cheng et al.,, 2016), instead, used only the estimated conversion rate as the score for each item, therefore, aiming to recommend the most relevant items to consumers. Netflix (Amatriain,, 2013) adopted a score function that linearly combines popularity and predicted rating for each video. Yang et al., (2019) proposes a linear programming based algorithm for online advertisement bid, without considering the ranking. As we will introduce more details in section˜2, Marketplace A also utilizes a score-based ranking method whose objective combines different sources of revenue (Ge et al.,, 2021). In the case of “cold start”, i.e., the launch of a new advertisement/item, where the data for conversation rate estimation is limited, Ye et al., (2023) proposed a bandit algorithm that significantly increased the success rate of cold start in field experiments.
@@
An inevitable drawback of the score-based method is that even though online platforms aim to achieve Pareto efficiency over multi-objective (Lin et al.,, 2019), it is hard to guarantee the performance of individual objectives. An alternative approach is to formulate the problem as a mathematical programming problem, that is, maximizing an objective function while setting constraints on some other metrics and resources. Such a problem can be formulated as a mixed integer programming (MIP) problem (Celis et al.,, 2017), and sometimes referred to as an advertisement allocation problem (Mehta et al.,, 2013). We can also associate this problem with some well-studied literature on assignment problem (Ross and Soland,, 1975) and knapsack problem (Salkin and De Kluyver,, 1975). However, as the number of constraints increases, the MIP also becomes increasingly harder to solve, requiring various approximation techniques. A more comprehensive review can be referred to (Cattrysse and Van Wassenhove,, 1992; Pentico,, 2007; Karp et al.,, 1990). Particularly, Celis et al., (2017) considered the ranking problem with fairness constraints. Based on LP relaxation, they recovered integral solutions with small constraint violations. Asadpour et al., (2023) studied a ranking model that aims to maximize consumer engagement for online retail, proposing an algorithm by LP relaxation and randomized rounding. Further, through LP-relaxation and primal-dual pairs, Google (Bhalgat et al.,, 2012) addressed an advertisement allocation problem with nested packing constraints, and then approximated it with an LP-based primal-dual algorithm; Microsoft (Chen et al.,, 2011) proposed a bidding algorithm for advertisement allocation with budget constraints. Inspired by (Chen et al.,, 2011), Alibaba (Zhang et al.,, 2018)) developed a similar algorithm for a global advertisement allocation problem that optimized over accumulated search queries.
@@
In the context of recommender systems, the primary design objective is to ensure low latency, that is, the ability to generate rankings within a strictly limited response time. Consequently, unlike assortment optimization problems (which we will discuss shortly), ranking problems typically do not treat item interactions as a central concern. Most studies therefore assume that position weights (often referred to as “position bias”) are independent of the item assortment (Chen et al.,, 2023; Agarwal et al.,, 2019; Hofmann et al.,, 2013) and develop methods such as propensity-score-based estimation to correct for this bias (Chen et al.,, 2023; Hofmann et al.,, 2013; Joachims et al.,, 2017).
@@
Our paper also relates closely to the literature that studies the selection of objectives and trade-offs between muti-objectives. As mentioned earlier, given the variety of objectives, different online platforms may prioritize them differently. Other than maximizing revenue, online marketplaces might also need to consider extra objectives such as consumer utility, ranking relevance (Ghose et al.,, 2014), algorithm fairness (Geyik et al.,, 2019), and content diversity (Chen et al.,, 2024). For instance, online video platforms such as YouTube or Netflix emphasized more on ranking relevance (Covington et al.,, 2016; Amatriain,, 2013), while online marketplaces such as Tabao cared more about revenue (Hu et al.,, 2018). Naturally, each online marketplace has multiple objectives incorporated into the optimization framework. Agarwal et al., (2011) jointly optimized for initial clicks and post-click downstream. Xiao et al., (2017) considered the fairness conflict between individuals and the entire group. Chen et al., (2024) recommended that social media platforms should focus on increasing content diversification specifically for active consumers. Contrary to industry assumptions, their research revealed that diversified recommendations would enhance content consumption diversity only among active consumers. Adomavicius and Kwon, (2014) optimized for the aggregated diversity of a group of recommendations. Sumida et al., (2021) characterized the trade-off between revenue and utility under MNL, by constructing the objective function as a weighted sum of revenue and utility. Louca et al., (2019) proposed a similar MNL model optimization framework, weighting the sum of revenue and relevance. Li et al., (2025) considered a fairness metric evaluated by the click-throughs of different Ads and customer types. They proposed a two-stage allocation method for online advertisement, which balanced revenue and algorithm fairness. However, since the objectives might not always align, Rodriguez et al., (2012) addressed this issue using constraint optimization. On the other hand, as multi-objective optimization often resulted in many nondominated solutions and low pressure towards the Pareto front, Li et al., (2016) developed a stochastic ranking algorithm to tackle this problem.
@@
The work is also related to the literature studying assortment optimization problems, which refer to the challenge of selecting a subset of products to offer from a larger set, in an online setting, where the goal is often to maximize revenue or customer satisfaction (Kök et al.,, 2009; Caro and Gallien,, 2007; Gallego et al.,, 2020). Differing from ranking problems, assortment problems focus on studying the interaction among listed items and do not usually study the effect of orders. In assortment literature, many previous works adopt parametric models to model the purchase probability of items given a certain set of items, such as the independent demand model (IDM) and basic attrition model (BAM) (Gallego et al.,, 2019). Particularly, IDM assumes that the probability of sale for each item is independent of others, thus the optimal assortment would be the one including all items with positive revenue. However, IDM might be pessimistic in a way that when items are running out of stock, recapturing any of those sales to alternate products would be infeasible since the demands are treated as independent (Gallego et al.,, 2015). Under BAM, on the other hand, for a given assortment, the choice probability of each item is normalized by the summation of the utility of all items in that assortment. The often-used multinomial choice model (MNL) is a special case of BAM (McFadden et al.,, 1973; Luce,, 2005). Under the MNL model with no constraint, Talluri and Van Ryzin, (2004); Gallego et al., (2004) characterized the revenue-ordered property of the optimal assortment. Further, for the robust assortment optimization under MNL, the optimal assortment is still revenue-ordered (Rusmevichientong and Topaloglu,, 2012). However, in contrast to IDM, BAM might be optimistic about recapturing the sale of alternative items due to the self-normalization procedure (Gallego et al.,, 2015). Feldman et al., (2022) also emperically showed that the MNL model performed worse during multi-purchase events. Recently, there has been some research studying the ranking of the assortment of products for online retailing. Ferreira et al., (2022) studied the ranking of the selected assortment display to consumers. The proposed online learning algorithms also for maximum consumer engagement, which balanced between learning speed and optimality gap. As classic choice models for online retailing only gave limited consideration to substitution patterns, Jiang et al., (2025) proposed a high-dimensional choice as an alternative. Asadpour et al., (2023) focused on ranking assortment in the form of a submodular maximization problem. They developed offline approximation algorithms that give the maximum probability of purchases. Most of the previous works in the assortment literature focused on developing theory guarantees. The only empirical study we found on online assortment problems is Feldman et al., (2022). They novelly implemented the algorithm proposed in Rusmevichientong et al., (2010) and presented an empirical study at Alibaba comparing the multinomial logit model versus industry-level machine learning models for assortment.
@@ 2.3 Score-Based Ranking
The parameter $w\in\mathbb{R}_{+}$ governs the relative importance of organic revenue versus ad revenue. It is optimized at the impression level based on factors including the predicted PTR, position weights, ad rates, and item prices. The method for optimizing $w$ is detailed in Ge et al., (2021). Notably, the selection of the scoring function and the calibration of $w$ has gone through extreme tuning at Marketplace over an extended period, and we believe their production model optimally harnesses the power of the score-based algorithm.
@@ 3.1 Mixed-Integer Programming (MIP) Formulation for the SLR Problem
We highlight several key assumptions in our model setup. First, we assume that the position weight $h_{i}$ is independent of the PTR. While this independence may not hold universally across all sponsored listings, the assumption is motivated by the need for low-latency response times. This simplification is commonly adopted in industry practice (Chen et al.,, 2023). Second, while our ranking stage does not explicitly account for interactions among items in a sponsored listing due to latency considerations, we address potential dependencies through PTR estimation. That is, the estimated PTR for each item can depend on the overall assortment of items in the sponsored listing. However, the position weight $h_{i}$ remains fixed and is not influenced by the assortment. This setup reflects a deliberate design choice: we aim to capture item interactions during prediction stage rather than the optimization stage. Third, our model allows for the possibility that a consumer purchases multiple items from a single sponsored listing. This flexibility is consistent with platform objectives, as multiple purchases do not negatively impact the platform, sellers, or consumers. In practice, our partner marketplace data show that, among sponsored listings that led to purchases, most involved only one item, and the maximum number of items purchased from a single listing was four.
@@ 3.4 Why the LP-Based Algorithm Outperforms the Score-Based Algorithm
Several clarifications are in order. First, both our proposed algorithm and the score-based algorithm incorporate slot-specific weights $h_{i}$. As described in Section 2.3, the parameter $w$ in the score-based method is optimized at the impression level, accounting for predicted PTR, position weights, ad rates, and item prices (see Ge et al., 2021 for details) [^4] Our modeling idea could, in principle, be expressed in a score-based framework as $\delta_{j}=r_{j}\times t_{j}\times p_{j}+r_{j}\times ad_{j}\times p_{j}+\mu\times r_{j}$, where $\mu$ is the dual variable associated with the relevance constraint. We nonetheless prefer the LP-based formulation because it naturally accommodates hard constraints on relevance, consistent with marketplace goals of safeguarding consumer experience and long-term performance, while still maximizing revenue subject to those constraints. Moreover, the LP-based formulation possesses the modeling flexibility to accommodate additional constraints and can be effectively applied in an online setting.. Second, in designing our field experiment, we deliberately adopted a conservative test strategy. Specifically, while $w$ in the score-based algorithm is optimized at the impression level, our $\lambda$ is applied globally across all impressions. To ensure a fair and conservative benchmark, we set $\lambda$ above the average relevance level achieved by the score-based algorithm in pre-experiment data. This design allows us to test whether our LP-based approach can improve both revenue and relevance through a more principled formulation of the tradeoff.
@@ 5.2.2 Item Characteristics Favored by the LP-based Algorithm
Results on item characteristics favored by the LP-based algorithm in the selection stage are reported in Table 4. Two patterns are noteworthy. First, by more effectively capturing the revenue–relevance tradeoff, both LP90 and LP95 identify items with higher expected revenue, encompassing both organic and advertising revenue, relative to the score-based benchmark in the item selection stage. Second, our LP-based algorithms favor more popular items. A potential concern is that such selection may exacerbate inequities in online marketplaces, as niche products face reduced opportunities for promotion. An advantage of our LP-based framework, however, is its flexibility: fairness constraints can be readily incorporated (e.g., Lu et al., 2025), and we provide additional extensions along these lines in Appendix A.5.
@@ 6 Conclusion
Several important avenues remain for future research. First, while our empirical evidence focuses on sponsored listings, the framework could be extended to KR settings where ranking plays a critical role, such as in our partner marketplace. A key distinction is that sponsored items are typically specified to consumers in KR, a challenge that would need to be addressed in such extensions. Second, following industry practice, we estimated position weights ($h_{i}$) and PTR separately and treated them as given when optimizing the ranking process. Developing efficient methods to jointly estimate $h_{i}$ and PTR represents an interesting and important next step. Finally, although we provide preliminary evidence on the long-term effects of our LP-based algorithm, richer field experiments specifically designed to capture sustained outcomes would yield deeper insights into its long-run impact (Zhang et al., 2020). In particular, future investigations could examine how the estimation of PTR should be adapted to capture consumers’ behavioral changes under the new ranking algorithm, as well as how advertisers adjust their bidding and advertising strategies in response.

### v4

@@ 1 Introduction
A widely adopted approach for the optimization stage in practice is the score-based ranking algorithm, which assigns each item a composite score based on predefined criteria and ranks items according to these scores. These algorithms are favored for their computational efficiency, typically producing rankings within 0.1 seconds to satisfy real-time latency requirements. The scoring criteria generally integrate multiple objectives, such as predicted consumer action rate, item relevance, and potential revenue generation. In most implementations, platforms employ a linear combination of these objectives to construct the score function that best aligns with their business priorities (liu2015convolutional; cheng2016wide; covington2016deep; okura2017embedding; zhou2018deep).
@@ 1.1 Literature reivew
Our work provides an approach for the ranking step in the recommendation system. In a recommendation system, the first step is to estimate the conversion rate of displaying the impression by machine learning algorithms, which has been quite mature and provide reasonable estimations in the online advertising industry (liu2015convolutional; cheng2016wide; covington2016deep; okura2017embedding; zhou2018deep). After estimating the conversion rates, online platforms are tasked with developing a ranking model that displays candidate items to consumers in a designated order, which is the focus of this paper. A majority of previous studies focus on the score-based ranking model (zehlike2022fairness; renda2003web; montague2001relevance; manmatha2001modeling), that is, calculating a score for each item through a customized scoring function, and then selecting the displayed items by ranking the scores. While the ranking algorithms are similar among various studies, the scoring function, however, might vary, depending on the objectives of each online platform. For instance, covington2016deep deployed a deep neural network model and used logistic regressions to generate scores for each video content on YouTube. Google Play (cheng2016wide), instead, used only the estimated conversion rate as the score for each item, therefore, aiming to recommend the most relevant items to consumers. Netflix (amatriain2013big) adopted a score function that linearly combines popularity and predicted rating for each video. yang2019bid proposes a linear programming based algorithm for online advertisement bid, without considering the ranking. As we will introduce more details in section˜2, Marketplace A also utilizes a score-based ranking method whose objective combines different sources of revenue (ge2020relevance). In the case of “cold start”, i.e., the launch of a new advertisement/item, where the data for conversation rate estimation is limited, ye2023cold proposed a bandit algorithm that significantly increased the success rate of cold start in field experiments.
@@
An inevitable drawback of the score-based method is that even though online platforms aim to achieve Pareto efficiency over multi-objective (lin2019pareto), it is hard to guarantee the performance of individual objectives. An alternative approach is to formulate the problem as a mathematical programming problem, that is, maximizing an objective function while setting constraints on some other metrics and resources. Such a problem can be formulated as a mixed integer programming (MIP) problem (celis2017ranking), and sometimes referred to as an advertisement allocation problem (mehta2013online). We can also associate this problem with some well-studied literature on assignment problem (ross1975branch) and knapsack problem (salkin1975knapsack). However, as the number of constraints increases, the MIP also becomes increasingly harder to solve, requiring various approximation techniques. A more comprehensive review can be referred to (cattrysse1992survey; pentico2007assignment; karp1990optimal). Particularly, celis2017ranking considered the ranking problem with fairness constraints. Based on LP relaxation, they recovered integral solutions with small constraint violations. asadpour2023sequential studied a ranking model that aims to maximize consumer engagement for online retail, proposing an algorithm by LP relaxation and randomized rounding. Further, through LP-relaxation and primal-dual pairs, Google (bhalgat2012online) addressed an advertisement allocation problem with nested packing constraints, and then approximated it with an LP-based primal-dual algorithm; Microsoft (chen2011real) proposed a bidding algorithm for advertisement allocation with budget constraints. Inspired by (chen2011real), Alibaba (zhang2018whole)) developed a similar algorithm for a global advertisement allocation problem that optimized over accumulated search queries.
@@
In the context of recommender systems, the primary design objective is to ensure low latency, that is, the ability to generate rankings within a strictly limited response time. Consequently, unlike assortment optimization problems (which we will discuss shortly), ranking problems typically do not treat item interactions as a central concern. Most studies therefore assume that position weights (often referred to as “position bias”) are independent of the item assortment (chen2023bias; agarwal2019general; hofmann2013reusing) and develop methods such as propensity-score-based estimation to correct for this bias (chen2023bias; hofmann2013reusing; joachims2017unbiased).
@@
Our paper also relates closely to the literature that studies the selection of objectives and trade-offs between muti-objectives. As mentioned earlier, given the variety of objectives, different online platforms may prioritize them differently. Other than maximizing revenue, online marketplaces might also need to consider extra objectives such as consumer utility, ranking relevance (ghose2014examining), algorithm fairness (geyik2019fairness), and content diversity (chen2024effects). For instance, online video platforms such as YouTube or Netflix emphasized more on ranking relevance (covington2016deep; amatriain2013big), while online marketplaces such as Tabao cared more about revenue (hu2018reinforcement). Naturally, each online marketplace has multiple objectives incorporated into the optimization framework. agarwal2011click jointly optimized for initial clicks and post-click downstream. xiao2017fairness considered the fairness conflict between individuals and the entire group. chen2024effects recommended that social media platforms should focus on increasing content diversification specifically for active consumers. Contrary to industry assumptions, their research revealed that diversified recommendations would enhance content consumption diversity only among active consumers. adomavicius2014optimization optimized for the aggregated diversity of a group of recommendations. sumida2021revenue characterized the trade-off between revenue and utility under MNL, by constructing the objective function as a weighted sum of revenue and utility. louca2019joint proposed a similar MNL model optimization framework, weighting the sum of revenue and relevance. li2021online considered a fairness metric evaluated by the click-throughs of different Ads and customer types. They proposed a two-stage allocation method for online advertisement, which balanced revenue and algorithm fairness. However, since the objectives might not always align, rodriguez2012multiple addressed this issue using constraint optimization. On the other hand, as multi-objective optimization often resulted in many nondominated solutions and low pressure towards the Pareto front, li2016stochastic developed a stochastic ranking algorithm to tackle this problem.
@@
The work is also related to the literature studying assortment optimization problems, which refer to the challenge of selecting a subset of products to offer from a larger set, in an online setting, where the goal is often to maximize revenue or customer satisfaction (kok2009assortment; caro2007dynamic; gallego2020approximation). Differing from ranking problems, assortment problems focus on studying the interaction among listed items and do not usually study the effect of orders. In assortment literature, many previous works adopt parametric models to model the purchase probability of items given a certain set of items, such as the independent demand model (IDM) and basic attrition model (BAM) (gallego2019revenue). Particularly, IDM assumes that the probability of sale for each item is independent of others, thus the optimal assortment would be the one including all items with positive revenue. However, IDM might be pessimistic in a way that when items are running out of stock, recapturing any of those sales to alternate products would be infeasible since the demands are treated as independent (gallego2015general). Under BAM, on the other hand, for a given assortment, the choice probability of each item is normalized by the summation of the utility of all items in that assortment. The often-used multinomial choice model (MNL) is a special case of BAM (mcfadden1973conditional; luce2005individual). Under the MNL model with no constraint, talluri2004revenue; gallego2004managing characterized the revenue-ordered property of the optimal assortment. Further, for the robust assortment optimization under MNL, the optimal assortment is still revenue-ordered (rusmevichientong2012robust). However, in contrast to IDM, BAM might be optimistic about recapturing the sale of alternative items due to the self-normalization procedure (gallego2015general). feldman2022customer also emperically showed that the MNL model performed worse during multi-purchase events. Recently, there has been some research studying the ranking of the assortment of products for online retailing. ferreira2022learning studied the ranking of the selected assortment display to consumers. The proposed online learning algorithms also for maximum consumer engagement, which balanced between learning speed and optimality gap. As classic choice models for online retailing only gave limited consideration to substitution patterns, jiang2020high proposed a high-dimensional choice as an alternative. asadpour2023sequential focused on ranking assortment in the form of a submodular maximization problem. They developed offline approximation algorithms that give the maximum probability of purchases. Most of the previous works in the assortment literature focused on developing theory guarantees. The only empirical study we found on online assortment problems is feldman2022customer. They novelly implemented the algorithm proposed in rusmevichientong2010dynamic and presented an empirical study at Alibaba comparing the multinomial logit model versus industry-level machine learning models for assortment.
@@ 2.3 Score-Based Ranking
The parameter $w\in\mathbb{R}_{+}$ governs the relative importance of organic revenue versus ad revenue. It is optimized at the impression level based on factors including the predicted PTR, position weights, ad rates, and item prices. The method for optimizing $w$ is detailed in ge2020relevance. Notably, the selection of the scoring function and the calibration of $w$ has gone through extreme tuning at Marketplace over an extended period, and we believe their production model optimally harnesses the power of the score-based algorithm.
@@ 3.1 Mixed-Integer Programming (MIP) Formulation for the SLR Problem
We highlight several key assumptions in our model setup. First, we assume that the position weight $h_{i}$ is independent of the PTR. While this independence may not hold universally across all sponsored listings, the assumption is motivated by the need for low-latency response times. This simplification is commonly adopted in industry practice (chen2023bias). Second, while our ranking stage does not explicitly account for interactions among items in a sponsored listing due to latency considerations, we address potential dependencies through PTR estimation. That is, the estimated PTR for each item can depend on the overall assortment of items in the sponsored listing. However, the position weight $h_{i}$ remains fixed and is not influenced by the assortment. This setup reflects a deliberate design choice: we aim to capture item interactions during prediction stage rather than the optimization stage. Third, our model allows for the possibility that a consumer purchases multiple items from a single sponsored listing. This flexibility is consistent with platform objectives, as multiple purchases do not negatively impact the platform, sellers, or consumers. In practice, our partner marketplace data show that, among sponsored listings that led to purchases, most involved only one item, and the maximum number of items purchased from a single listing was four.
@@ 3.4 Why the LP-Based Algorithm Outperforms the Score-Based Algorithm
Several clarifications are in order. First, both our proposed algorithm and the score-based algorithm incorporate slot-specific weights $h_{i}$. As described in Section 2.3, the parameter $w$ in the score-based method is optimized at the impression level, accounting for predicted PTR, position weights, ad rates, and item prices (see ge2020relevance for details) [^4] Our modeling idea could, in principle, be expressed in a score-based framework as $\delta_{j}=r_{j}\times t_{j}\times p_{j}+r_{j}\times ad_{j}\times p_{j}+\mu\times r_{j}$, where $\mu$ is the dual variable associated with the relevance constraint. We nonetheless prefer the LP-based formulation because it naturally accommodates hard constraints on relevance, consistent with marketplace goals of safeguarding consumer experience and long-term performance, while still maximizing revenue subject to those constraints. Moreover, the LP-based formulation possesses the modeling flexibility to accommodate additional constraints and can be effectively applied in an online setting.. Second, in designing our field experiment, we deliberately adopted a conservative test strategy. Specifically, while $w$ in the score-based algorithm is optimized at the impression level, our $\lambda$ is applied globally across all impressions. To ensure a fair and conservative benchmark, we set $\lambda$ above the average relevance level achieved by the score-based algorithm in pre-experiment data. This design allows us to test whether our LP-based approach can improve both revenue and relevance through a more principled formulation of the tradeoff.
@@ 5.2.2 Item Characteristics Favored by the LP-based Algorithm
Results on item characteristics favored by the LP-based algorithm in the selection stage are reported in Table 4. Two patterns are noteworthy. First, by more effectively capturing the revenue–relevance tradeoff, both LP90 and LP95 identify items with higher expected revenue, encompassing both organic and advertising revenue, relative to the score-based benchmark in the item selection stage. Second, our LP-based algorithms favor more popular items. A potential concern is that such selection may exacerbate inequities in online marketplaces, as niche products face reduced opportunities for promotion. An advantage of our LP-based framework, however, is its flexibility: fairness constraints can be readily incorporated (e.g., luzhu2025), and we provide additional extensions along these lines in Appendix A.5.
@@ 6 Conclusion
Several important avenues remain for future research. First, while our empirical evidence focuses on sponsored listings, the framework could be extended to KR settings where ranking plays a critical role, such as in our partner marketplace. A key distinction is that sponsored items are typically specified to consumers in KR, a challenge that would need to be addressed in such extensions. Second, following industry practice, we estimated position weights ($h_{i}$) and PTR separately and treated them as given when optimizing the ranking process. Developing efficient methods to jointly estimate $h_{i}$ and PTR represents an interesting and important next step. Finally, although we provide preliminary evidence on the long-term effects of our LP-based algorithm, richer field experiments specifically designed to capture sustained outcomes would yield deeper insights into its long-run impact (zhanglongterm2020). In particular, future investigations could examine how the estimation of PTR should be adapted to capture consumers’ behavioral changes under the new ranking algorithm, as well as how advertisers adjust their bidding and advertising strategies in response.
Online Appendix